the length of two pieces of wire are in the ratio 4:7.if the lenght of the longer piece is 3.5m, what is the length of the shorter piece?
Answer by ptaylor(2198)

Let x=length of the shorter piece:
Now we are basically told that 4 is to 7 as x is to 3.5m and this is written as follows:
4/7=x/3.5 Now we cross multiply: (x)(7)=(4)(3.5)
7x=12.5
x=1.786m


4/7=x/3.5 Now lets look at this another way: We'll multiply both sides by (7)(3.5), the LCM, to clear the fractions and we get:
4(3.5)(7/7)=x(7)(3.5/3.5)
12.5=7x-----------------------we get the same result


Hope this helps----ptaylor